### v2

# Apparent Selection Pressure for Channel Capacity in Bacterial Chemotactic Sensors

Abstract
Bacterial chemotactic sensing converts noisy chemical signals into running and tumbling. We analyze the static sensing limits of mixed Tar/Tsr chemoreceptor clusters in individual Escherichia coli cells using a heterogeneous Monod–Wyman–Changeux (MWC) model. By sweeping a seven-dimensional parameter space, we compute three sensing performance metrics—channel capacity, effective Hill coefficient, and dynamic range. Across E. coli–like parameter regimes, we consistently observe pronounced local maxima of channel capacity, whereas neither the effective Hill coefficient nor the dynamic range exhibit comparable optimization. The capacity-achieving input distribution is bimodal, which implies that individual cells maximize information by sampling both low- and high concentration regimes. Together, these results suggest that, at the individual-cell level, channel capacity may be selected for in E. coli receptor clusters.
I Introduction
Sensing is fundamental to life. Across all domains of biology, sensing is a noisy conversation between the environment and an organism’s decision-making. With bad sensors can come bad decisions, and so evolution might have driven organisms to maximize sensors’ information processing capabilities for survival.
A long-standing biophysical hypothesis thus theorizes that evolution pushes sensory pathways to maximize information transfer from environment to sensor output. For instance, measurements indicate that fruit fly embryos transmit information on gene regulation from chemical gradients approximately $90\%$ of the theoretical maximum [18]. Nicotinic acetylcholine receptors in the uncoupled limit appear to have maximized their channel capacity [11], which is correlated with maximal cooperativity and maximal dynamic range for Monod-Wyman-Changeux molecules [9]. However, potassium voltage-gated channels within animals have been found to operate far from an information-maximizing regime and have evolved toward lower information transfer [6]. Similarly, single mammalian cells only transmit sufficient information for binary decision-making, and additional information is limited by an upstream bottleneck [5]. Finally, evolutionary channels can only transmit a few bits of information [16].
A model organism of study for this are bacteria. Their bacterial chemotactic receptors, responsible for the movement of bacteria cells in response to environmental chemical stimuli, have been the subject of a number of beautiful models [11, 19, 13, 14]. In order to survive, E. coli cells must take in information about chemoattractant concentration [2] so as to move to maximize chemoattractant intake. Recent measurements in E. coli cells have demonstrated that they operate close to optimal information transfer for gradient-climbing [12]. A recent study showed that bacteria optimize predictive information transfer given resource constraints [17]. Furthermore, basic theoretical analyses say that populations of bacteria should maximize predictive information subject to constraints on resources [10].
However, no studies have addressed the channel capacities of individual bacterium. While some studies on E. coli chemotactic receptors focus on the population level and dynamical behavior of cells [9, 10], the present study provides an analysis of the chemotactic ability of a single E. coli cell. Specifically, the mixed Tar/Tsr chemoreceptor clusters within the E. coli cell are modeled with a heterogeneous Monod-Wyman-Changeux molecule [11] under the assumption of adaptation to negligible background chemoattractant concentration [4]. Three sensor performance metrics (channel capacity, effective Hill coefficient, and dynamic range) are obtained by sweeping a seven-dimensional parameter space. By locating literature-based strains in these seven-dimensional landscapes, we find that channel capacity appears to have been driven to a local maximum for all available strains. Dynamic range and effective Hill coefficient are positively correlated with channel capacity across the seven-dimensional parameter space, but upon closer examination, do not appear to have been driven to local maxima in the seven-dimensional parameter space by evolution, although dynamic range is near the maximum of what it could be. This implies that channel capacity is the subject of evolutionary selective pressure, adding to the literature on optimized properties of the bacterial chemotactic sensors.
II Background
II.1 Bacterial Chemotaxis
Motile bacteria regulate their motion in response to spatial variations in chemical concentrations so that, on average, trajectories are biased up attractant gradients and down repellent gradients. This bias does not require the cell to know the global gradient. Instead, as the cell swims, it continuously samples the local environment and adjusts its behavior based on changes in concentration it has recently experienced. Bacterial chemotaxis is characterized by an alternating sequence of runs and tumbles [11]. During a run, several flagella rotate in the same direction, form a bundle, and propel the cell along an approximately straight path. A tumble occurs when the cell reorients itself by breaking the bundle and rotating its flagella in different directions (Fig. 1).
Figure 1: Schematic demonstrating the run–tumble behavior of E. coli.
The chemical signals that bias this run–tumble behavior are called chemoattractants or chemorepellents, which bind to chemoreceptors embedded in the cell membrane. In E. coli, the dominant chemoreceptor types are Tar and Tsr. Tar primarily recognizes aspartate and its analogues such as MeAsp, whereas Tsr primarily recognizes serine. At the cell pole, many Tar and Tsr receptors organize with the adaptor CheW and the histidine kinase CheA in a signaling cluster whose members switch conformations collectively. The cluster alternates between an inactive state, in which CheA kinase activity is suppressed, and an active state, in which CheA autophosphorylation from ATP is stimulated [11]. Mixed Tar/Tsr composition tunes the cluster’s sensitivity to distinct ligands, and the cooperative coupling ensures that modest changes in ligand occupancy can produce large changes in the cluster’s output.
The downstream link to motility is mediated by the CheY response regulator. Phosphorylated CheA transfers its phosphate to CheY, producing CheY-P. CheY-P binds to the flagellar motor and increases its clockwise rotation, causing a tumble. In contrast, when the CheA output is reduced, CheY-P levels drop and the flagellar motor rotates counterclockwise to induce a run. For attractants such as MeAsp or serine, ligand binding stabilizes the inactive cluster state and thereby lowers CheA/CheY-P signaling, biasing the motor toward longer runs; decreases in attractant or the presence of repellents shifts the balance toward the active state and increases the likelihood of reorientation. As a result, bacterial chemotaxis produces net motion toward higher concentrations of attractants and away from higher concentrations of repellents (Fig. 2).
Figure 2: Schematic demonstrating bacterial chemotactic behavior and receptor response. (a) In a spatially uniform environment (no attractant or repellent), the run–tumble trajectory is an unbiased random walk with no net drift. (b) Positive chemotaxis in an attractant gradient: run segments are longer when the cell swims up the gradient and shorter when it swims down, yielding a net displacement toward higher concentration. Bar indicates increasing concentration. (c) Negative chemotaxis in a repellent gradient: more frequent reorientation when moving up-gradient biases motion toward lower concentration.
II.2 Monod-Wyman-Changeux Model of Bacterial Chemotactic Receptors
Consider a receptor cluster composed of $N_{1}$ Tar receptors (class $i=1$) and $N_{2}$ Tsr receptors (class $i=2$). The cluster switches collectively between two conformations $s\in\{I,A\}$ (inactive $I$, active $A$). At an external ligand concentration $c$, binding to receptor class $i$ in conformation $s$ is characterized by the state-dependent dissociation constant $K_{d}^{(s),i}$. At any instant, some integer number of Tar receptors and Tsr receptors are bound to ligand. Let $\varepsilon_{s}$ denote the conformational free energy of state $s$ and $\beta=1/(k_{B}T)$.
From Ref. [11], we have
[…]
Define the allosteric constant $L_{0}$, which represents the ratio of the inactive and active state statistical weights in the absence of ligand, given as
$L_{0}\equiv e^{-\beta(\varepsilon_{I}-\varepsilon_{A})}.$ (2)
By substituting Eq. 2 into 1 and dividing the numerator and denominator by $e^{-\beta\varepsilon_{A}}\!\left(1+\frac{c}{K_{d}^{(A),1}}\right)^{N_{1}}\!\left(1+\frac{c}{K_{d}^{(A),2}}\right)^{N_{2}}$ yields
$p(c)=\frac{1}{1+L_{0}\left(\frac{1+c/K_{d}^{(I),1}}{1+c/K_{d}^{(A),1}}\right)^{N_{1}}\left(\frac{1+c/K_{d}^{(I),2}}{1+c/K_{d}^{(A),2}}\right)^{N_{2}}}.$ (3)
For attractants with $K_{d}^{(I),i}<K_{d}^{(A),i}$, $p(c)$ decreases monotonically with $c$ (Fig. 3 with parameters from Table LABEL:tab:1). Thus, the receptor cluster is more likely to be inactive at attractant concentrations, which allows the E. coli to have a run-biased trajectory up the attractant gradient.
Adaptation [19] is ignored not only for simplicity, but because in natural environments with large bacterial populations, chemoattractant and chemoattractant gradients constantly vanish [4]. Thus, negligible background concentration is assumed.
Figure 3: Probability of the active state of a Tar/Tsr receptor cluster for the attractant MeAsp for ten different strains. Normalized activity decreases monotonically with MeAsp concentration on a log scale because ligand stabilizes the inactive state.
Table 1: Strain-specific sensor parameters [14].
Strain $N_{1}$ $N_{2}$ $L_{0}$ $K_{d}^{(I),1}$ $K_{d}^{(A),1}$ $K_{d}^{(I),2}$ $K_{d}^{(A),2}$
1 4.95 16.5 $3.499\times 10^{-4}$ 0.0492 0.1096 0.0345 0.1099
2 4.00 8.0 $1.672\times 10^{-2}$ 0.0492 0.1096 0.0345 0.1099
3 4.39 4.39 $7.331\times 10^{-2}$ 0.0492 0.1096 0.0345 0.1099
4 18.7 6.24 $1.705\times 10^{-3}$ 0.0492 0.1096 0.0345 0.1099
5 14.0 0.00 $6.676\times 10^{-2}$ 0.0492 0.1096 0.0345 0.1099
6 29.8 0.00 $2.534\times 10^{-3}$ 0.0492 0.1096 0.0345 0.1099
7 73.5 0.00 $2.985\times 10^{-7}$ 0.0492 0.1096 0.0345 0.1099
8 0.00 9.85 $1.722\times 10^{-2}$ 0.0492 0.1096 0.0345 0.1099
9 0.00 15.2 $1.709\times 10^{-3}$ 0.0492 0.1096 0.0345 0.1099
10 0.00 32.3 $1.062\times 10^{-6}$ 0.0492 0.1096 0.0345 0.1099
II.3 Channel capacity
Channel capacity is the maximum mutual information that the receptor channel can transmit between the input ligand concentration and the output (activity). Mutual information $I(C;S)$ quantifies the average amount of information obtained about the input $c$ by observing the output $s\in\{I,A\}$. For each $c$, the conditional law is given by Eq. 3: $p(s{=}1\mid c)=p_{\mathrm{active}}(c)$ and $p(s{=}0\mid c)=1-p_{\mathrm{active}}(c)$. Let $p(c)$ be an arbitrary input distribution over concentrations and define the output marginal $p(s)=\int p(c)\,p(s\mid c)\,dc$. The mutual information in bits is given as
$I(C;S)=\sum_{s\in\{0,1\}}\int p(c)\,p(s\mid c)\,\log_{2}\!\frac{p(s\mid c)}{p(s)}\,dc.$ (4)
Since the channel has a binary output, $I(C;S)$ is between 0 and $1$ bits. $I(C;S)=0$ bits indicates that $c$ and $s$ are independent, while $I[C;S]=1$ bit indicates that the output is deterministic. Noise in the sensor or environment decreases mutual information. Channel capacity is the maximum mutual information obtained by choosing the most informative input distribution. Mathematically, it is the supremum over input distributions, given as
$C=\sup_{p(c)}I(C;S).$ (5)
Biologically, one must argue that the sensor changes its input concentration via movement so as to saturate its information processing capabilities.
III Methods
To evaluate how well the receptor cluster can sense ligand concentration, its sensor properties can be described by three performance metrics: Channel capacity, effective hill coefficient, and dynamic range. We compute these metrics from the MWC activity curve $p_{\mathrm{active}}(c)$ and sweep seven model parameters: the allosteric constant $L_{0}$, the four state-dependent dissociation constants $K_{d}^{(I),1}$, $K_{d}^{(A),1}$, $K_{d}^{(I),2}$, $K_{d}^{(A),2}$, and the Tar/Tsr receptor copy numbers $N_{1}$ and $N_{2}$.
To visualize how these metrics depend on parameters, we plot two-dimensional slices through the seven-dimensional landscape as heatmaps. In each panel, the two plotted parameters span the axes, while the remaining five parameters are aggregated by taking the per-pixel median over all sampled values, yielding a representative summary of performance across the full sweep. Axes for $K_{d}^{(s),i}$ and $L_{0}$ are displayed on logarithmic scales, while axes for $N_{1}$ and $N_{2}$ are linear.
For visualization purposes, heatmap color scales are rescaled using percentile clipping, with limits set by the 2nd and 98th percentiles of the metric values within each panel. This procedure prevents extreme outliers from saturating the colormap and allows internal structure near biologically relevant regions to be resolved. All numerical analyses, correlations, and gradient calculations are performed on the unclipped data.
III.1 Channel capacity
Channel capacity $C$ is numerically calculated via the Blahut–Arimoto algorithm [3, 1], which iteratively reweights the input distribution to monotonically increase $I(C;S)$ until convergence. In addition to the capacity value $C^{*}$, the algorithm also outputs the optimal Blahut–Arimoto input distribution $p^{*}(c)$ and the optimal output occupancy $p^{*}(s)$. $p^{*}(c)$ is the distribution of ligand concentrations that would realize that capacity, and $p^{*}(s)$ shows the fraction of time that the receptor ends up inactive versus active states at capacity.
We model receptor-cluster sensing as a discrete memoryless channel whose input is ligand concentration and whose output is the binary activity state of the cluster. The concentration axis is represented by a log-spaced grid $\{c_{j}\}_{j=1}^{M}$ of $M$ bins chosen to sweep over concentrations that showcase both the plateaus and the steep changes in activity curves of Fig. 3. Refining the log-spaced grid $\{c_{j}\}$ left qualitative conclusions unchanged. For each bin $c_{j}$, the output $s\in\{0,1\}$ (0 == inactive, 1 == active) is distributed according to the MWC response (Eq. 3):
$P(s=1\mid c_{j})$ $=p_{\mathrm{active}}(c_{j}),$ (6)
$P(s=0\mid c_{j})$ $=1-p_{\mathrm{active}}(c_{j}).$
Let $p(c_{j})$ denote the prior probability that the input lies in bin $j$. The induced output marginal is $p(s)=\sum_{j=1}^{M}p(c_{j})\,P(s\mid c_{j})$ for $s\in\{0,1\}$.
The mutual information (in bits) between the input concentration $C$ and the output state $S$ on this grid is
$I(C;S)=\sum_{j=1}^{M}p(c_{j})\sum_{s\in\{0,1\}}P(s\mid c_{j})\,\log_{2}\!\frac{P(s\mid c_{j})}{p(s)}.$ (7)
The static channel capacity is the maximum of $I(C;S)$ over all input priors on the grid,
$C=\max_{\,p(c)}I(C;S),$ (8)
and because the output is binary one always has $0\leq C\leq 1$ bit per independent observation.
To compute $C$ and a maximizing prior, we use the Blahut–Arimoto (BA) algorithm for fixed discrete channels. Starting from any initial prior $p^{(0)}(c_{j})$ with $\sum_{j}p^{(0)}(c_{j})=1$, define at iteration $t$ the output marginal
$p^{(t)}(s)=\sum_{j=1}^{M}p^{(t)}(c_{j})\,P(s\mid c_{j}),\qquad s\in\{0,1\}.$ (9)
For compactness, define the per-input “row score”
$\ell^{(t)}_{j}=\sum_{s\in\{0,1\}}P(s\mid c_{j})\,\ln\!\frac{P(s\mid c_{j})}{\,p^{(t)}(s)\,}$, (10)
The prior is then updated by normalized reweighting:
$p^{(t+1)}(c_{j})=\frac{\exp\!\big(\ell^{(t)}_{j}\big)}{\displaystyle\sum_{k=1}^{M}\exp\!\big(\ell^{(t)}_{k}\big)},\qquad j=1,\dots,M,$ (11)
which guarantees $\sum_{j}p^{(t+1)}(c_{j})=1$. These iterations monotonically increase $I(C;S)$ and converge to a capacity-achieving prior $p^{*}(c_{j})$. We report the capacity $C^{*}=I_{p^{*}}(C;S)$ (converted to bits via $\log_{2}$), together with the optimal input distribution $p^{*}(c_{j})$ and the induced optimal output occupancy
$p^{*}(s)=\sum_{j=1}^{M}p^{*}(c_{j})\,P(s\mid c_{j}),\qquad s\in\{0,1\}.$ (12)
Because the output is binary, the maximizing prior typically concentrates probability near concentrations whose outputs are most distinct.
III.2 Dynamic range
Dynamic range is defined as the change in mean activity between the basal (no input) and saturated (very high input) states. We have
$p_{0}$ $=$ $p(0)=\frac{1}{1+L_{0}}$ (13)
and
$p_{\infty}$ $=$ $\lim_{c\to\infty}p(c)=\frac{1}{1+L_{0}\left(\frac{K_{d}^{(A),1}}{K_{d}^{(I),1}}\right)^{N_{1}}\left(\frac{K_{d}^{(A),2}}{K_{d}^{(I),2}}\right)^{N_{2}}}$ (14)
with
$\mathrm{DR}=\bigl|\,p_{\infty}-p_{0}\,\bigr|.$ (15)
III.3 Effective Hill coefficient
Cooperative ligand-receptor binding interactions can be fit to the Hill equation:
$p(c)=\frac{(c/K_{d})^{n_{\mathrm{eff}}}}{1+(c/K_{d})^{n_{\mathrm{eff}}}}.$ (16)
Physically, the effective Hill coefficient $n_{\mathrm{eff}}$ quantifies how abruptly the output changes as $c$ crosses its midpoint, indicating that the binding sites act in concert (cooperativity). Negative cooperativity comes from $|n_{\mathrm{eff}}|<1$, while positive cooperativity comes from $|n_{\mathrm{eff}}|>1$. Greater deviations of $|n_{\mathrm{eff}}|$ from $1$ means a more switch-like response (small input changes cause large output changes near the concentration that leads to middling $p(c)$), but may narrow the dynamic range of the sensor.
To find the effective Hill coefficient, define the transition output level as the midpoint
$p^{*}\equiv\frac{p_{0}+p_{\infty}}{2},$ (17)
and define $c^{*}$ as the unique concentration that satisfy
$p(c^{*})=p^{*},$ (18)
and define the normalized activity
$a(c)\equiv\frac{p(c)-p_{0}}{p_{\infty}-p_{0}}.$ (19)
The effective Hill coefficient is [11]
$n_{\mathrm{eff}}=2\,\left.\frac{d}{d\ln c}\ln a(c)\right|_{c=c^{*}}.$ (20)
As shown in Supplementary Information,
$n_{\mathrm{eff}}$ $=$  $-\,4\,\frac{c^{*}\,p^{*}(1-p^{*})}{\,p_{\infty}-p_{0}\,}\,\Bigg[N_{2}\!\left(\frac{1}{K_{d}^{(I),1}+c^{*}}-\frac{1}{K_{d}^{(A),1}+c^{*}}\right)$ (21)
$+N_{2}\!\left(\frac{1}{K_{d}^{(I),2}+c^{*}}-\frac{1}{K_{d}^{(A),2}+c^{*}}\right)\Bigg].$
IV Results
The static sensing performance of the mixed Tar/Tsr receptor cluster is evaluated by varying seven parameters: $L_{0}$, $K_{d}^{(I),1}$, $K_{d}^{(A),1}$, $K_{d}^{(I),2}$, $K_{d}^{(A),2}$, $N_{1}$, $N_{2}$. For each set of parameters, three performance metrics are computed to evaluate the sensor’s response: dynamic range $\mathrm{DR}$, effective hill coefficient $n_{\mathrm{eff}}$, and channel capacity $C$. Two-dimensional slices of this seven-dimensional parameter space are visualized in heatmaps, with color scales indicating the magnitude of each metric. To anchor the analysis in biological data, the receptor composition data of ten strains of E. coli are extracted from literature and overlayed on the heatmap as white symbols (Table 1). These strain markers provide reference points for where real bacteria sit in parameter space.
To quantify relationships among the three static metrics across the 7D sweep, we computed pairwise correlation coefficient $R$ (Table 2).
All three metrics are positively correlated as was true for populations of MWC receptors [9]. Channel capacity correlates strongly with dynamic range ($R=0.978$) and moderately with the effective Hill coefficient ($R=0.600$); effective Hill coefficient and dynamic range are also moderately correlated ($R=0.481$). These correlations support the idea that high information transfer typically requires both a sufficiently steep response and a broad input span. Despite the strong correlations, we claim that it is possible to distinguish which quantities are potential selection pressures in a manner described below.
$C$ $n_{\mathrm{eff}}$ $\mathrm{DR}$
$C$ 1 0.600 0.978
$n_{\mathrm{eff}}$ 0.600 1 0.481
$\mathrm{DR}$ 0.978 0.481 1
Table 2: Correlation coefficient between performance metrics.
For clarity, this section displays a representative subset of heatmaps with strain reference points overlaid. See the supplementary information section for the full set of heatmaps across ten strains.
IV.1 Channel capacity
For all strains, the capacity-achieving input distribution $p^{*}(c)$ is strongly bi-modal, concentrating probability near the two endpoints of the response and showing a pronounced minimum at the midpoint. See Fig. 4. For a noisy binary receptor, observations are most informative when the output is near-deterministic ($p_{\mathrm{active}}(c)\approx 0$ or $\approx 1$) and least informative at the midpoint where the Bernoulli variance p(1p)p(1-p) is maximal. In other words, the information-maximizing strategy is spend time in both low and high concentration regions rather than to dwell at the middle. The corresponding optimal output distribution $p^{*}(s)$ is slightly biased toward the active state, which indicate mild channel asymmetry.
All optimal input distributions are similar despite differences in strain parameters. Figure 5 thus forms a testable prediction for these strains. Actual input distributions in natural environments in which gradients constantly vanish [4] might resemble the optimal input distributions shown here, with mutual informations close to capacity.
Figure 4: Optimal use of the bacterial chemotactic receptors for all strains using parameters in Table 1. At top, capacity-achieving input distribution $p^{*}(c)$ versus ligand concentration $c$ as obtained by the Blahut-Arimoto algorithm. At bottom, optimal output distribution $p^{*}(s)$ among the active and inactive state for all strains.
Figure 5: Channel capacity versus $K_{d}^{(A),1}$ for each of the ten strains holding all other parameters fixed.
Strain $C$ for strain $_{max} CC$ for strain / $C_{max}$
Strain 1 0.83 0.95 0.87
Strain 2 0.76 0.95 0.80
Strain 3 0.64 0.95 0.68
Strain 4 0.81 0.95 0.86
Strain 5 0.65 0.95 0.68
Strain 6 0.80 0.95 0.85
Strain 7 0.87 0.95 0.92
Strain 8 0.77 0.95 0.81
Strain 9 0.82 0.95 0.86
Strain 10 0.86 0.95 0.90
Table 3: Channel capacity by strain. With $C_{max}$ the maximum channel capacity found over the 7D parameter sweep, the channel capacity for each strain. All capacities are in bits.
Strain Gradient $\|\nabla C\|_{2}$ for strain Global max $\|\nabla C\|_{2}$ Gradient norm ratio
Strain 1 0.02 5.75 0.00
Strain 2 0.10 5.75 0.02
Strain 3 0.47 5.75 0.08
Strain 4 0.06 5.75 0.01
Strain 5 0.62 5.75 0.11
Strain 6 0.12 5.75 0.02
Strain 7 0.14 5.75 0.02
Strain 8 0.40 5.75 0.07
Strain 9 0.15 5.75 0.03
Strain 10 0.11 5.75 0.02
Table 4: Local flatness of channel capacity as measured by the L2 norm of the gradient of channel capacity. Gradient L2 norm for each strain versus the global maximum of the gradient L2 norm over the seven-dimensional parameter sweep. All gradients estimated numerically.
Figure 5 maps channel capacity $C$ across a key MWC parameter. A binary receptor conveys information only if there exist input concentrations whose outputs are meaningfully different, which is determined by the dynamic range. When the two endpoints are well separated, the Blahut–Arimoto algorithm concentrates input probability near those regions, making the outputs easy to distinguish and yielding a high $C$; when the dynamic range is small, all inputs appear similar at the output and $C$ collapses. Because channel capacity is strongly correlated with dynamic range ($R=0.978$), the dominant qualitative parameter dependencies of $C$ can be understood directly from the dynamic range expression, and we therefore defer a detailed discussion of these trends to the following subsection.
Fig. 5 suggests that strains are poised at a local maximum of channel capacity in seven-dimensional parameter space. This is confirmed by numerical calculation of gradient norms of channel capacity versus the maximum possible gradient norm of channel capacity in Table 4. However, we are at a local maximum rather than a global maximum of channel capacity based on Table 3.
IV.2 Dynamic range
Dynamic range quantifies how far the receptor output can move between the basal and saturated limits. From Eq. (15) and as visually corroborated by the heatmaps, dynamic range is governed by at least three qualitative trends. First, $\mathrm{DR}$ is suppressed when $K_{d}^{(A),i}\approx K_{d}^{(I),i}$, because $K_{d}^{(A),i}/K_{d}^{(I),i}\approx 1$ and $p_{\infty}\approx p_{0}$. Second, in the attractant regime $K_{d}^{(I),i}<K_{d}^{(A),i}$, increasing the separation between active- and inactive-state affinities and/or increasing receptor copy numbers $(N_{1},N_{2})$ amplifies the product term and pushes $p_{\infty}$ farther from $p_{0}$, thereby increasing $\mathrm{DR}$ toward its upper bound. Third, $\mathrm{DR}$ can remain small even when $K_{d}^{(A),i}\neq K_{d}^{(I),i}$ due to extreme allosteric bias. For $L_{0}\ll 1$, $p_{0}\approx 1$ and typically $p_{\infty}\approx 1$ unless the product term is large enough to overcome the small prefactor $L_{0}$; for $L_{0}\gg 1$, $p_{0}\approx 0$ and typically $p_{\infty}\approx 0$ as well.
Since dynamic range and channel capacity are strongly correlated (Table 2), channel capacity can be observed following the same qualitative trends.
Dynamic range looks, based on representative heatmaps, to be maximized in parameter space in Fig. 6. Like for channel capacity, it looks like strains sit on the edge of a plateau for dynamic range. And according to Table 5, this dynamic range maximization is global, with most strain’s dynamic ranges near the global maximum of $1$. However, Table 6 tells a slightly different story: Numerical gradients of dynamic range are very large for many of the strains, so the heatmaps are actually misleading upon closer inspection. From this, we conclude that dynamic range is large and but potentially not maximized for these strains.
Figure 6: A representative heatmap suggesting that dynamic range is also at a global maximum in parameter space, though see Table 6. Dynamic range versus $K_{d}^{(A),1}$ and $K_{d}^{(A),2}$ appears to be locally maximized by strain parameters for strain 11.
Strain $\mathrm{DR}$ for strain $\mathrm{DR}_{max}$ $\mathrm{DR}$ for strain / $\mathrm{DR}_{max}$
Strain 1 1.00 1.00 1.00
Strain 2 0.98 1.00 0.98
Strain 3 0.93 1.00 0.93
Strain 4 1.00 1.00 1.00
Strain 5 0.94 1.00 0.94
Strain 6 1.00 1.00 1.00
Strain 7 1.00 1.00 1.00
Strain 8 0.98 1.00 0.98
Strain 9 1.00 1.00 1.00
Strain 10 1.00 1.00 1.00
Table 5: Dynamic range by strain compared with sweep maxima. $\mathrm{DR}_{max}$ is the maximum dynamic range over the 7D sweep.
Strain Gradient $\|\nabla\mathrm{DR}\|_{2}$ for strain Global max of $\|\nabla\mathrm{DR}\|_{2}$ (7D) Gradient norm ratio
Strain 1 0.27 6.00 0.05
Strain 2 0.29 6.00 0.05
Strain 3 0.47 6.00 0.08
Strain 4 0.69 6.00 0.12
Strain 5 2.71 6.00 0.45
Strain 6 2.98 6.00 0.50
Strain 7 3.00 6.00 0.50
Strain 8 1.62 6.00 0.27
Strain 9 1.87 6.00 0.31
Strain 10 2.33 6.00 0.39
Table 6: Lack of flatness of dynamic range as measured by the L2 norm of the gradient of dynamic range. Gradient L2 norm for each strain versus the global maximum of the gradient L2 norm over the seven-dimensional parameter sweep. All gradients estimated numerically.
IV.3 Effective Hill coefficient
Effective Hill coefficient quantifies how sharply the receptor output switches between its basal and saturated limits around the midpoint concentration. From Eq. (21) (and consistent with the heatmaps), the effective Hill coefficient is governed by several qualitative trends. First, $|n_{\mathrm{eff}}|$ increases with receptor copy number: because $N_{1}$ and $N_{2}$ enter linearly in the sensitivity term, larger Tar/Tsr cluster sizes yield a more switch-like transition near the midpoint. Second, $|n_{\mathrm{eff}}|$ increases as the difference between inactive- and active-state dissociation constants increases. When $K_{d}^{(I),i}\approx K_{d}^{(A),i}$ the difference terms $\left(\frac{1}{K_{d}^{(I),i}+c^{*}}-\frac{1}{K_{d}^{(A),i}+c^{*}}\right)$ vanish, producing a shallow response and thus small $|n_{\mathrm{eff}}|$; conversely, larger $K_{d}^{(A),i}/K_{d}^{(I),i}$ amplifies these terms and steepens the curve. Third, the prefactor $4\,\frac{c^{*}\,p^{*}(1-p^{*})}{p_{\infty}-p_{0}}$ shows that endpoint saturation matters: if the activity curve does not meaningfully span distinct low- and high-concentration plateaus ($p_{\infty}\approx p_{0}$), then the midpoint-based steepness becomes ill-conditioned, and $n_{\mathrm{eff}}$ can be undefined when the response is effectively flat.
Based on Fig. 7, it is not locally maximized for strain 1; and see Supplementary Information for other strains. This intuition is borne out by Table 7, which shows that effective Hill coefficients are far from a global maximum, and Table 8, which shows that strains have large Hill coefficient gradient norms. From this, we conclude that effective Hill coefficient is not locally or globally maximized.
Figure 7: A heatmap suggesting that effective Hill coefficient is not at a local maximum in parameter space. Effective Hill coefficient versus $K_{d}^{(A),1}$ and $K_{d}^{(A),2}$ is locally maximized by strain parameters for strain 11.
Strain $n_{\mathrm{eff}}$ at bio dot Max $n_{\mathrm{eff}}$ over 7D sweep Ratio of $n_{\mathrm{eff}}$
Strain 1 5.03 176.96 0.03
Strain 2 2.74 176.96 0.02
Strain 3 1.87 176.96 0.01
Strain 4 4.48 176.96 0.03
Strain 5 2.23 176.96 0.01
Strain 6 4.46 176.96 0.03
Strain 7 11.10 176.96 0.06
Strain 8 2.63 176.96 0.01
Strain 9 3.94 176.96 0.02
Strain 10 8.43 176.96 0.05
Table 7: Effective Hill coefficient by strain compared with sweep maxima.
Strain Gradient $\|\nabla n_{\mathrm{eff}}\|_{2}$ at bio point Global max $\|\nabla n_{\mathrm{eff}}\|_{2}$ (7D) Ratio of gradient norms
Strain 1 9.54 903.03 0.01
Strain 2 3.83 903.03 0.00
Strain 3 2.29 903.03 0.00
Strain 4 9.10 903.03 0.01
Strain 5 3.86 903.03 0.00
Strain 6 10.58 903.03 0.01
Strain 7 30.48 903.03 0.03
Strain 8 4.53 903.03 0.01
Strain 9 8.29 903.03 0.01
Strain 10 21.81 903.03 0.02
Table 8: Local flatness of the effective Hill coefficient: gradient norm at each strain versus the global maximum over the 7D sweep.
V Conclusions
This study examined the static sensing limits of mixed Tar/Tsr chemoreceptor clusters in E. coli using a heterogeneous MWC framework. By sweeping a seven-parameter space and computing three performance metrics—channel capacity, effective Hill coefficient, and dynamic range—we mapped how receptor composition and state-dependent affinities shape the information available at the receptor output level. While dynamic range tends to be large, it does not seem to be maximized based on gradient analyses, and the effective Hill coefficient appears to not be maximized either. Channel capacity reveals itself as locally maximized by all E. coli strains. This is revealed even though all three characterizations of performance are correlated, as in Ref. [9].
Thus, E. coli, like nACh receptors [11] and Drosophila embryos [18], appear to maximize information transfer in a quantitative way. From channel capacity calculations, E. coli is able to maximize information transfer by spending time in both low and high concentration regions. This makes testable predictions for what kinds of input bacteria receive in harsh, natural environments. This is hard to measure, but based on the analyses so far, it appears that channel capacity has been selected for in E. coli.
This lends another information-theoretic analysis to the long list of information-theoretic and optimality analyses [12, 17, 4]. Other work has shown that efficient prediction as in Ref. [17] also explains cultured neurons spiking [8], human behavior [7], and salamander retinal ganglion neuron spiking [15]. It is possible that all these optimality principles are required to explain E. coli, or that high performance on one optimality principle is correlated with high performance on another. Only by testing in other organisms will we find out which optimality principles generalize to other evolved organisms and which do not.